OpenStudy (anonymous):

approximate these coordinates to the nearest hundreth y=x^2+2 y=x+3

OpenStudy (anonymous):

what are you axactly asking for

OpenStudy (anonymous):

just to approximate these coordinates to the nearest hundredth and if there is no solution state it

OpenStudy (anonymous):

he wants to know that \[x = .5(1-\sqrt{5}), y = .5(7-\sqrt{5})\] or \[x = .5(1+\sqrt{5}), y = .5(7+\sqrt{5})\]